### v3

@@ 5 Literature Overview
Repeatedly applying RIIA we can remove irrelevant alternatives [^17] An alternative $a\in A\in\mathcal{X}$ is irrelevant with respect to $A$ if $p(a,A)=0$. from A to reach the menu $A^{{}^{\prime}}=\{x:x\in A,p(x,A)>0\}$. Note that $p(x,A)=p(x,A^{{}^{\prime}})$ for all $x\in A^{{}^{\prime}}$.
@@
$A\cap\bar{E}\neq\phi$ for some $A\in\mathcal{X}$ implies $p(a^{*},A)=0$.
Proof. Suppose not. Then there exists $A\in\mathcal{X}$ such that $A\cap\bar{E}\neq\phi$ and $p(a^{*},A)>0$. For $x\in A\cap\bar{E}$, we have by REG that $p(a^{*},\{x\})>0$ or $p(x,\{x\})\neq 1$. Since $x\in\bar{E}$, $p(x,B)=1$ for some $B\in\mathcal{X}$. Applying REG leads to $p(x,\{x\})=1$ which is a contradiction.
@@
Claim 4: For all $A\in\mathcal{X}$ and $a,b\in A$: $p(a,A)>0$, $p(b,\{b\})\neq 1$ and $p(a,A\setminus\{b\})>p(a,A)$ imply $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$ for all $a\in A^{*}$.
@@
A stochastic choice function $p:X^{*}\times\mathcal{X}$ has a Random Utility Model (RUM) representation if there exists a distribution $\nu$ over the set of all strict total orders on $X^{*}$, $\mathcal{L}(X^{*})$ such that:

### v4

@@ 5 Literature Overview
Repeatedly applying RIIA we can remove irrelevant alternatives [^17] An alternative $a\in A\in\mathcal{X}\setminus\phi$ is irrelevant with respect to $A$ if $p(a,A)=0$. from A to reach the menu $A^{{}^{\prime}}=\{x:x\in A,p(x,A)>0\}$. Note that $p(x,A)=p(x,A^{{}^{\prime}})$ for all $x\in A^{{}^{\prime}}$.
@@
$A\cap\bar{E}\neq\phi$ for some $A\in\mathcal{X}\setminus\phi$ implies $p(a^{*},A)=0$.
Proof. Suppose not. Then there exists $A\in\mathcal{X}\setminus\phi$ such that $A\cap\bar{E}\neq\phi$ and $p(a^{*},A)>0$. For $x\in A\cap\bar{E}$, we have by REG that $p(a^{*},\{x\})>0$ or $p(x,\{x\})\neq 1$. Since $x\in\bar{E}$, $p(x,B)=1$ for some $B\in\mathcal{X}\setminus\phi$. Applying REG leads to $p(x,\{x\})=1$ which is a contradiction.
@@
Claim 4: For all $A\in\mathcal{X}\setminus\phi$ and $a,b\in A$: $p(a,A)>0$, $p(b,\{b\})\neq 1$ and $p(a,A\setminus\{b\})>p(a,A)$ imply $\frac{p(a,A\setminus\{b\})}{p(a,A)}=\frac{1}{1-p(b,\{b\})}$ for all $a\in A^{*}$.
@@
A stochastic choice function $p:X^{*}\times\mathcal{X}\setminus\phi$ has a Random Utility Model (RUM) representation if there exists a distribution $\nu$ over the set of all strict total orders on $X^{*}$, $\mathcal{L}(X^{*})$ such that: